If a tenant has a percentage lease at 5% of sales what is the annual rent from a tenant if
monthly sales are $35 per SF?

If the apartment building has 100 units, 100,000 SF and 250 parking spaces what is the parking ratio for the building?
Answer #1

1. Apartment has 100 units with total area of 100,000 Square Feet (SF). Thus each unit is of 1000 SF.

Monthly sales = 1000 *35= 35000

Annual Rent amount = 35000*5%*12 (months) = USD 21000.

2. Parking Ratio = Parking space/No. of Units = 250/100 = 2.5.